The Ka of a monoprotic weak acid is 3.84 × 10?-3. What is the percent ionization of a 0.132 M solution of this acid?   

I keep getting negative numbers inside the square root of the quadratic, and I'm not sure what I'm doing wrong.
Answer #1

HA dissociates as:

HA          ----->     H+   + A-
0.132                 0         0
0.132-x               x         x


Ka = [H+][A-]/[HA]
Ka = x*x/(c-x)
Assuming x can be ignored as compared to c
So, above expression becomes

Ka = x*x/(c)
so, x = sqrt (Ka*c)
x = sqrt ((3.84*10^-3)*0.132) = 2.251*10^-2

since x is comparable c, our assumption is not correct
we need to solve this using Quadratic equation
Ka = x*x/(c-x)
3.84*10^-3 = x^2/(0.132-x)
5.069*10^-4 - 3.84*10^-3 *x = x^2
x^2 + 3.84*10^-3 *x-5.069*10^-4 = 0
This is quadratic equation (ax^2+bx+c=0)
a = 1
b = 3.84*10^-3
c = -5.069*10^-4

Roots can be found by
x = {-b + sqrt(b^2-4*a*c)}/2a
x = {-b - sqrt(b^2-4*a*c)}/2a

b^2-4*a*c = 2.042*10^-3

roots are :
x = 2.068*10^-2 and x = -2.452*10^-2

since x can't be negative, the possible value of x is
x = 2.068*10^-2

% dissociation = (x*100)/c
= 2.068*10^-2*100/0.132
= 15.7 %
Answer: 15.7 %